If sulfur dioxide diffuses through a porous barrier at the rate of 2.4 mol/min, at what rate would chlorine diffuse? Assume that the gases behave ideally.
1
Expert's answer
2019-10-17T07:18:54-0400

Molecular mass of "SO_2(m_1)=64 gm"

Molecular mass of "Cl_2(m_2)=35.5 \times 2=71 \ gm"

By using Graham law of diffusion,

"\frac{r_1}{r_2}=\sqrt{\frac{m_2}{m_1}}"

where "r_1" and "r_2" are rate of diffusion respectively.

"r_1=2.4 \ mol/min"

Using the above equation,

"\frac{2.4}{r_2}=\sqrt{\frac{71}{64}}\implies r_2=2.28 \ mol/min"